# Approximate independence of permutation mixtures

### v2

@@ 2.4 Method of cumulants
$b_{2\ell+1}\sim 2\left(\frac{2}{\pi}\right)^{2\ell+1}\cdot(2\ell+1)!\qquad\text{as }\ell\to\infty.$
@@
$\kappa_{(1,2\ell+1,0,\ldots,0)}=a_{2\ell+1}\sim(-1)^{\ell+1}\frac{\pi}{n-1}\left(\frac{2\mu}{\pi}\right)^{2\ell+2}\cdot(2\ell+1)!\qquad\text{as}\ell\to\infty.$
@@ 3 Upper bound via the doubly centered expansion
In this section we prove the first part (4) of Theorem 1.2, by developing an orthogonal expansion of the likelihood ratio in terms of a set of “doubly centered” functions. In Section 3.1 we first provide a direct proof of the previous toy example and discuss the intuitions, and then present the proof for the general case in Section 3.2. One key step is a new upper bound of elementary symmetric polynomials for a sequence summing into zero, which we present in Section 3.3.
@@ 3.1 A warm-up example
Similar to Section 2, this section focuses on the toy Gaussian model in (1) as a warm-up example. The proof relies critically on the following representation of the Gaussian likelihood ratio:
$\frac{\varphi(x-\theta)}{\varphi(x)}=\exp\left(\theta x-\frac{\theta^{2}}{2}\right)=\cosh(\mu x)\exp\left(-\frac{\mu^{2}}{2}\right)+\sinh(\mu x)\frac{\theta}{\mu}\exp\left(-\frac{\mu^{2}}{2}\right),\quad\theta\in\{\pm\mu\},$ (14)
@@
=(14)${=}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\left(1+\tanh(\mu X_{i})\frac{\theta_{\pi(i)}}{\mu}\right)\right]$
@@
where in (a), the contribution of $S=\varnothing$ cancels with $-1$, and $g_{1}=0$ is used for $|S|=1$. This proves (1).
@@
By contrast, the above approach works directly with the ratio $\varphi(x-\theta)/\varphi_{0}(x)=1+\Psi(x,\theta)$, where $\Psi(x,\theta):=\tanh(\mu x)\frac{\theta}{\mu}$. This representation has two benefits with respect to an expansion in Hermite polynomials. First, the product structure of $\mathbb{Q}_{n}$ guarantees that for any $\theta_{i}\in\{\pm\mu\}$, the functions $\{\Pi_{i\in S}\Psi(X_{i},\theta_{i}):S\subseteq[n]\}$ are automatically orthogonal in $L^{2}(\mathbb{Q}_{n})$, so no change of measure is required. Second, and more subtly, for any $x\in\mathbb{R}^{d}$, the function $\theta\mapsto\Psi(x,\theta)$ is automatically mean-zero when $\theta\sim\mathrm{Unif}(\{\pm\mu\})$. This property guarantees that the first-order term in (3.1) vanishes, which is crucial to obtaining the correct dependence on $\mu$, and indirectly ensures that the coefficients $g_{\ell}$ are typically small. On the other hand, as Section 2 makes clear, large values of the moments $m_{\alpha}$ or cumulants $\kappa_{\alpha}$ lead to the failure of the moment and cumulant methods. Further comparisons between the above two approaches are discussed in Section 5.4.
Several challenges arise when attempting to implement this idea in general. There will typically not be a simple expression for the analogue of $\Psi(x,\theta)$ in the general case, and the coefficients no longer admit an explicit formula such as (18). We address these challenges in Section 3.2.
@@ 3.2 Proof of the general case
This section is devoted to the proof of the first part (4) of Theorem 1.2. Motivated by (14) and (15) in the warm-up example, we write
@@
As we shall see, the first of these equations guarantees that products functions in $\{\Psi_{i}\}_{i\in[n]}$ are orthogonal in $L^{2}(\mathbb{Q}_{n})$, whereas the second guarantees that $\mathbb{E}_{I\sim\mathrm{Unif}[n]}\Psi_{I}(\mathrm{d}x)=0$. Based on (19) and (20), we express the likelihood ratio as
@@
In the special case $\ell=n$, Lemma 3.1 coincides with the Hadamard-type inequality for permanents [CLL06], even without the condition that all row sums are zero. However, for $\ell<n$, this condition becomes crucial: without this condition, if $A$ is the all-one matrix, the LHS of Lemma 3.1 would be a much larger quantity $\binom{n}{\ell}$. The proof of Lemma 3.1 turns out to be involved and is the central theme of Section 3.3.
Applying Lemma 3.1 to (23), for a given subset $S$ with $|S|=\ell\geq 2$ we get
@@
In addition, for $\ell=1$ the LHS of Lemma 3.1 is zero since all row sums are zero. Therefore, (23) gives that
@@
which is the first part (4) of Theorem 1.2.
@@ 3.3 A key inequality
This section is devoted to the proof of Lemma 3.1, which consists of several steps. First we show that it suffices to prove Lemma 3.1 with identical rows. Let $r_{1},\cdots,r_{\ell}\in{\mathcal{H}}$ be the rows of $A$, where =${\mathcal{H}}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}=0\}$ is a Hilbert space equipped with the Euclidean inner product. Writing the LHS of Lemma 3.1 as a function of $(r_{1},\dots,r_{\ell})$, it is clear that
@@
Let $L(x_{1},\cdots,x_{n})$ be a symmetric multilinear form from a Hilbert space $({\mathcal{H}},|\cdot|)$ to either $\mathbb{R}$ or $\mathbb{C}$. Suppose that
@@
To apply Lemma 3.2, we note that the multilinear form $L$ in Lemma 3.1 is clearly symmetric, and Lemma 3.1 precisely asks for an upper bound on the operator norm of $L$. By Lemma 3.2, it suffices to establish Lemma 3.1 when all rows of $A$ are identical, denoted by a vector $x\in{\mathcal{H}}$; also, by scaling we may assume that $\sum_{i=1}^{n}x_{i}^{2}=n$. In this case, the target quantity in Lemma 3.1 becomes $e_{\ell}(x):=\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}x_{i}$, the elementary symmetric polynomial of $x=(x_{1},\cdots,x_{n})$. The central inequality of this section is summarized in the following theorem.
@@
    If $x\in\mathbb{C}^{n}$ with $\sum_{i=1}^{n}x_{i}=0$ and $\sum_{i=1}^{n}|x_{i}|^{2}=n$, then (define $10^{0}:=$ 1)
@@
It is clear that the inequality (25) in the real case proves Lemma 3.1 after the reduction through Lemma 3.2. The complex inequality (24) will also be useful in Section 4. Again, we remark that the condition $\sum_{i=1}^{n}x_{i}=0$ is crucial for Theorem 3.3, for $x=(1,\ldots,1)$ would lead to $e_{\ell}(x)=\binom{n}{\ell}$. Theorem 3.3 gives a bound on $|e_{\ell}(x)|$ when $e_{1}(x)$ and $e_{2}(x)$ are known. It therefore adds to a recent line of work [GY14, MRT19, DHH20, Tao23] proving bounds on elementary symmetric polynomials given the values of two consecutive lower-order elementary symmetric polynomials. For real $x\in\mathbb{R}^{n}$, this prior work can be used to deduce the bound
@@
However, the application to our problem requires to have the best possible constant $A=1$ (note that the example in (18) implies that $A<1$ is impossible). For example, the quantity $f(\mu)$ in (2) could be arbitrarily close to $1$, and only $A=1$ makes the geometric sum $\sum_{\ell}A^{\ell}f(\mu)^{\ell}$ converge; we will also see similar scenarios in the application of Lemma 4.5 in Section 4. As shown in Theorem 3.3, this constant turns out to be achievable (possibly at some tolerable price of $\mathrm{poly}(\ell)$) using the saddle point method.
Proof of Theorem 3.3 (First Part).
@@
by choosing $r^{2}=\frac{\ell}{n-\ell}$ for $11\leq\ell\leq n-$ 1, $r\to 0$ for $\ell=0$, and $r\to\infty$ for $\ell=n$, establishing the first inequality of (24). The second step of (24) simply follows from Stirling’s approximation. ∎
The proof of the second part (25) of Theorem 3.3 is more involved. First, relying critically on a property of real-rooted polynomials (which fails for the complex case), we argue that the coordinates of the real maximizer $x$ of $|e_{\ell}(x)|$ can only take two values. The evaluation of $|e_{\ell}(x)|$ in the resulting simplified scenario is still highly challenging, and we apply a more careful saddle point analysis to arrive at (25). We defer the details to Appendix B.
@@ 4 Upper bound via matrix permanents
The bounds presented in Section 3 are based on the identity
@@
As mentioned in Section 1.2, this expression may be viewed as an expansion of the matrix permanent in (9) around the all-ones matrix. The benefit of this approach is that each of the terms in this expansion are orthogonal in $L^{2}(\mathbb{Q}_{n})$. However, our proof relies crucially on Lemma 3.2, which is used to crudely upper bound each permanent appearing in the above expansion by the permanent of a matrix with identical rows. This bound therefore ignores any additional structure in the likelihood ratio, and fails to give tight results when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})>1$.
In this section, we derive refined bounds for ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})>1$ by treating the $\chi^{2}$ divergence between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ directly. This gives proofs of the other upper bounds (5) and (6) in Theorem 1.2. We also show how to use similar ideas to obtain Theorem 1.6. In Section 4.1 we express the $\chi^{2}$ divergence between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ as a proper matrix permanent, and upper bound it in two ways, described in Section 4.2 and Section 4.3 respectively, both of which rely on the use of complex random variables.
@@ 4.1 Divergence as a matrix permanent
We first recall the expression given in Section 1.2 for $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ as a matrix permanent.
@@
By Lemma 4.1 and Lemma 4.2, the computation of the $\chi^{2}$ divergence $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ reduces to upper bounding the permanent of $A$. Moreover, in light of the van der Waerden conjecture/Egorychev–Falikman theorem [vdW26, Ego81, Fal81], which states that $\mathrm{Perm}(A)\geq\frac{n!}{n^{n}}$ for all $n\times n$ doubly stochastic matrices $A$, showing that $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=O(1)$ amounts to showing that $\mathrm{Perm}(A)$ is nearly as small as is possible for a doubly-stochastic matrix. Motivated by the eigenstructure of $A$ established in Lemma 4.2, we write $A=UDU^{\top}$ for an orthogonal matrix $U\in\mathbb{R}^{n\times n}$ and a diagonal matrix $D=\mathrm{diag}(\lambda_{1},\cdots,\lambda_{n})$ consisting of the eigenvalues. By a generalized Cauchy–Binet formula for permanents [MM65, Theorem 2.1], it holds that
@@
To relate the permanent view (27) with the doubly centered expansion in Section 3, we refer to Appendix A for some identities between $S_{\ell}$ and related quantities. In particular, we note that the permanent $\mathrm{Perm}(U_{n-\ell,\ell,0,\ldots,0})$ is exactly a multiple of $e_{\ell}(u_{2})$, the elementary symmetric polynomial of second column vector $u_{2}$ of $U$. Therefore, Theorem 3.3 could be used to upper bound this quantity. In addition, Lemma 3.2 reduces the general quantity $\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})$ to the case $\mathrm{Perm}(U_{n-\ell,\ell,0,\ldots,0})$ with identical columns, so the arguments in Section 3 leads to an upper bound of (27) in terms of $\sum_{i=2}^{n}\lambda_{i}\leq{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ by Lemma 4.2. However, this approach does not capture additional structures of $(\lambda_{2},\dots,\lambda_{n})$ such as the spectral gap. To this end, it might be a natural idea to establish pointwise upper bounds on $\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})$. We have not succeeded in this approach and leave it as an open direction; instead, we choose to upper bound the individual sum $S_{\ell}$ or the entire sum $\sum_{\ell=0}^{n}S_{\ell}$. These form the topics of the subsequent sections.
4.2 Proof of Theorem 1.2 via bounding the entire sum
In this section, we upper bound the entire sum $\sum_{\ell=0}^{n}S_{\ell}$ in (27) and use it to prove Theorem 1.2. The central result of this section is the following upper bound:
@@
Comparing this expression with (27), Lemma 4.3 shows that replacing all coefficients in (27) by $1$ gives an upper bound. The proof of Lemma 4.3 relies on the use of complex normal random variables. Recall that $z\sim{\mathcal{C}}{\mathcal{N}}(0,1)$ denotes $z=x+\mathrm{i}y$ with independent $x,y\sim{\mathcal{N}}(0,\frac{1}{2})$, and $z\sim{\mathcal{C}}{\mathcal{N}}(0,\Sigma)$ denotes a vector $z=x+\mathrm{i}y$ with independent $x,y\sim{\mathcal{N}}(0,\frac{\Sigma}{2})$ for real PSD matrices $\Sigma$. It is known that the moments of $z\sim{\mathcal{C}}{\mathcal{N}}(0,1)$ are
$\mathbb{E}[z^{m}\bar{z}^{n}]=\begin{cases}n!&\text{if }m=n,\\ 0&\text{if }m\neq n.\end{cases}$ (28)
@@
Lemma 4.4 follows from the classical Isserlis’ theorem [Iss18] (or Wick’s formula) for products of joint normal random variables, and we include a proof in Section C.7 for completeness. Now we present the proof of Lemma 4.3.
Proof of Lemma 4.3.
Let $A=UDU^{\top}$ be the matrix in Lemma 4.1, and $P=UD^{1/2}$. Then
@@
=$\overset{\rm(a)}{=}\frac{1}{n^{n}}\sum_{\begin{subarray}{c}(\ell_{1},\ldots,\ell_{n})\in\mathbb{N}^{n}\\ \sum_{i=1}^{n}\ell_{i}=n\end{subarray}}\binom{n}{\ell_{1},\ldots,\ell_{n}}\lambda_{1}^{\ell_{1}}\cdots\lambda_{n}^{\ell_{n}}\mathbb{E}\left[|z_{1}|^{2\ell_{1}}\cdots|z_{n}|^{2\ell_{n}}\right]$
=$\overset{\rm(b)}{=}\frac{n!}{n^{n}}\sum_{\begin{subarray}{c}(\ell_{1},\ldots,\ell_{n})\in\mathbb{N}^{n}\\ \sum_{i=1}^{n}\ell_{i}=n\end{subarray}}\lambda_{1}^{\ell_{1}}\cdots\lambda_{n}^{\ell_{n}},$
@@
Next we prove the second upper bound (5) of Theorem 1.2 from Lemma 4.3. Based on (27) and Lemma 4.3, we have
Thanks to the spectral gap bound in Lemma 4.2, we have $\lambda_{i}\in\left[0,1-\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\right]$, so that
$\overset{\rm(c)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\cdot\frac{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})-1}\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\overset{\rm(d)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\left(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\right),$
where (c) follows from Lemma 4.2, (d) uses $\log x\leq x-1$ for all $x>0$. For the inequality (6), we have
$\overset{\rm(e)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\cdot\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1}{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}\log({\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1)\overset{\rm(f)}{\leq}({\mathsf{C}}_{\chi^{2}}({\mathcal{P}})+1)\log({\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1),$
where (e) is due to $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\leq 1+{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$ and
and (f) uses ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$. Therefore the upper bounds (5) and (6) are established.
4.3 Proof of Theorem 1.6 via bounding the individual sum
In this section, we upper bound the individual sum $S_{\ell}$ in (27) and use it to prove Theorem 1.6. The central result of this section is the following upper bound:

### v3

@@ 3.4 Method of cumulants
[…]
@@
[…]
@@ 4 Upper bound via the doubly centered expansion
In this section we prove the first part (4) of Theorem˜1.2, by developing an orthogonal expansion of the likelihood ratio in terms of a set of “doubly centered” functions. In Section˜4.1 we first provide a direct proof of the previous toy example and discuss the intuitions, and then present the proof for the general case in Section˜4.2. One key step is a new upper bound of elementary symmetric polynomials for a sequence summing into zero, which we present in Section˜4.3.
@@ 4.1 A warm-up example
Similar to Section˜3, this section focuses on the toy Gaussian model in (1) as a warm-up example. The proof relies critically on the following representation of the Gaussian likelihood ratio: for $\theta\in\{\pm\mu\}$,
$\frac{\varphi(x-\theta)}{\varphi(x)}=\exp\left(\theta x-\frac{\theta^{2}}{2}\right)=\cosh(\mu x)\exp\left(-\frac{\mu^{2}}{2}\right)+\sinh(\mu x)\frac{\theta}{\mu}\exp\left(-\frac{\mu^{2}}{2}\right),$ (16)
@@
[…]
@@
where in (a), the contribution of $S=\varnothing$ cancels with 1-1, and $g_{1}=0$ is used for $|S|=1$. This proves (1).
@@
By contrast, the above approach works directly with the ratio $\varphi(x-\theta)/\varphi_{0}(x)=1+\Psi(x,\theta)$, where $\Psi(x,\theta):=\tanh(\mu x)\frac{\theta}{\mu}$. This representation has two benefits with respect to an expansion in Hermite polynomials. First, the product structure of $\mathbb{Q}_{n}$ guarantees that for any $\theta_{i}\in\{\pm\mu\}$, the functions $\{\Pi_{i\in S}\Psi(X_{i},\theta_{i}):S\subseteq[n]\}$ are automatically orthogonal in $L^{2}(\mathbb{Q}_{n})$, so no change of measure is required. Second, and more subtly, for any $x\in\mathbb{R}^{d}$, the function $\theta\mapsto\Psi(x,\theta)$ is automatically mean-zero when $\theta\sim\mathrm{Unif}(\{\pm\mu\})$. This property guarantees that the first-order term in (4.1) vanishes, which is crucial to obtaining the correct dependence on $\mu$, and indirectly ensures that the coefficients $g_{\ell}$ are typically small. On the other hand, as Section˜3 makes clear, large values of the moments $m_{\alpha}$ or cumulants $\kappa_{\alpha}$ lead to the failure of the moment and cumulant methods. Further comparisons between the above two approaches are discussed in Section˜6.3.
Several challenges arise when attempting to implement this idea in general. There will typically not be a simple expression for the analogue of $\Psi(x,\theta)$ in the general case, and the coefficients no longer admit an explicit formula such as (20). We address these challenges in Section˜4.2.
@@ 4.2 Proof of the general case
This section is devoted to the proof of the first part (4) of Theorem˜1.2. Motivated by (16) and (17) in the warm-up example, we write
@@
As we shall see, the first of these equations guarantees that products of functions in $\{\Psi_{i}\}_{i\in[n]}$ are orthogonal in $L^{2}(\mathbb{Q}_{n})$, whereas the second guarantees that $\mathbb{E}_{I\sim\mathrm{Unif}[n]}\Psi_{I}(\mathrm{d}x)=0$. Based on (21) and (22), we express the likelihood ratio as
@@
In the special case $\ell=n$, Lemma˜4.1 coincides with the Hadamard-type inequality for permanents [CLL06], even without the condition that all row sums are zero. However, for $\ell<n$, this condition becomes crucial: without this condition, if $A$ is the all-ones matrix, the LHS of Lemma˜4.1 would be a much larger quantity $\binom{n}{\ell}$. The proof of Lemma˜4.1 turns out to be involved and is the central theme of Section˜4.3.
Applying Lemma˜4.1 to (25), for a given subset $S$ with $|S|=\ell\geq 2$ we get
@@
In addition, for $\ell=1$ the LHS of Lemma˜4.1 is zero since all row sums are zero. Therefore, (25) gives that
@@
which is the first part (4) of Theorem˜1.2.
@@ 4.3 A key inequality
This section is devoted to the proof of Lemma˜4.1, which consists of several steps. First we show that it suffices to prove Lemma˜4.1 with identical rows. Let $r_{1},\cdots,r_{\ell}\in{\mathcal{H}}$ be the rows of $A$, where ${\mathcal{H}}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}=0\}$ is a Hilbert space equipped with the Euclidean inner product. Writing the LHS of Lemma˜4.1 as a function of $(r_{1},\dots,r_{\ell})$, it is clear that
@@
Let $L(x_{1},\cdots,x_{n})$ be a symmetric multilinear form from a Hilbert space $({\mathcal{H}},\langle\cdot,\cdot\rangle)$ to either $\mathbb{R}$ or $\mathbb{C}$. Suppose that
@@
To apply Lemma˜4.2, we note that the multilinear form $L$ in Lemma˜4.1 is clearly symmetric, and Lemma˜4.1 precisely asks for an upper bound on the operator norm of $L$. By Lemma˜4.2, it suffices to establish Lemma˜4.1 when all rows of $A$ are identical, denoted by a vector $x\in{\mathcal{H}}$; also, by scaling we may assume that $\sum_{i=1}^{n}x_{i}^{2}=n$. In this case, the target quantity in Lemma˜4.1 becomes $e_{\ell}(x):=\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}x_{i}$, the elementary symmetric polynomial of $x=(x_{1},\cdots,x_{n})$. The central inequality of this section is summarized in the following theorem.
@@
    If $x\in\mathbb{C}^{n}$ with $\sum_{i=1}^{n}x_{i}=0$ and $\sum_{i=1}^{n}|x_{i}|^{2}=n$, then (define $0^{0}:=1$)
@@
It is clear that the inequality (27) in the real case proves Lemma˜4.1 after the reduction through Lemma˜4.2. The complex inequality (26) will also be useful in Section˜5. Again, we remark that the condition $\sum_{i=1}^{n}x_{i}=0$ is crucial for Theorem˜4.3, for $x=(1,\ldots,1)$ would lead to $e_{\ell}(x)=\binom{n}{\ell}$. Theorem˜4.3 gives a bound on $|e_{\ell}(x)|$ when $e_{1}(x)$ and $e_{2}(x)$ are known. It therefore adds to a recent line of work [GY14, MRT19, DHH20, Tao23] proving bounds on elementary symmetric polynomials given the values of two consecutive lower-order elementary symmetric polynomials. For real $x\in\mathbb{R}^{n}$, this prior work can be used to deduce the bound
@@
However, the application to our problem requires to have the best possible constant $A=1$ (note that the example in (20) implies that $A<1$ is impossible). For example, the quantity $f(\mu)$ in (2) could be arbitrarily close to $1$, and only $A=1$ makes the geometric sum $\sum_{\ell}A^{\ell}f(\mu)^{\ell}$ converge; we will also see similar scenarios in the application of Lemma˜5.5 in Section˜5. As shown in Theorem˜4.3, this constant turns out to be achievable (possibly at some tolerable price of $\mathrm{poly}(\ell)$) using the saddle point method.
Proof of Theorem˜4.3 (First Part).
@@
by choosing $r^{2}=\frac{\ell}{n-\ell}$ for $1\leq\ell\leq n-1$, $r\to 0$ for $\ell=0$, and $r\to\infty$ for $\ell=n$, establishing the first inequality of (26). The second step of (26) simply follows from Stirling’s approximation. ∎
After the initial draft of this paper, we were informed that the same programs in Lemma˜4.2 and the first part of Theorem˜4.3 have already been employed in [Roo15]. However, the second part (27) of Theorem˜4.3 is still new, with a more involved proof. First, relying critically on a property of real-rooted polynomials (which fails for the complex case), we argue that the coordinates of the real maximizer $x$ of $|e_{\ell}(x)|$ can only take two values. The evaluation of $|e_{\ell}(x)|$ in the resulting simplified scenario is still challenging, and we apply a more careful saddle point analysis to arrive at (27). We defer the details to Appendix˜B.
@@ 5 Upper bound via matrix permanents
The bounds presented in Section˜4 are based on the identity
@@
As mentioned in Section˜1.2, this expression may be viewed as an expansion of the matrix permanent in (7) around the all-ones matrix. The benefit of this approach is that each of the terms in this expansion are orthogonal in $L^{2}(\mathbb{Q}_{n})$. However, our proof relies crucially on Lemma˜4.2, which is used to crudely upper bound each permanent appearing in the above expansion by the permanent of a matrix with identical rows. This bound therefore ignores any additional structure in the likelihood ratio, and fails to give tight results when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})>1$.
In this section, we derive refined bounds for ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})>1$ by treating the $\chi^{2}$ divergence between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ directly. This gives proofs of the other upper bounds (5) and (6) in Theorem˜1.2. We also show how to use similar ideas to obtain Theorem˜1.5. In Section˜5.1 we express the $\chi^{2}$ divergence between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ as a proper matrix permanent, and upper bound it in two ways, described in Section˜5.2 and Section˜5.3 respectively, both of which rely on the use of complex random variables.
@@ 5.1 Divergence as a matrix permanent
We first recall the expression given in Section˜1.2 for $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ as a matrix permanent.
@@
By Lemma˜5.1 and Lemma˜5.2, the computation of the $\chi^{2}$ divergence $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ reduces to upper bounding the permanent of $A$. Moreover, in light of the van der Waerden conjecture/Egorychev–Falikman theorem [vdW26, Ego81, Fal81], which states that $\mathrm{Perm}(A)\geq\frac{n!}{n^{n}}$ for all $n\times n$ doubly stochastic matrices $A$, showing that $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=O(1)$ amounts to showing that $\mathrm{Perm}(A)$ is nearly as small as is possible for a doubly-stochastic matrix. Motivated by the eigenstructure of $A$ established in Lemma˜5.2, we write $A=UDU^{\top}$ for an orthogonal matrix $U\in\mathbb{R}^{n\times n}$ and a diagonal matrix $D=\mathrm{diag}(\lambda_{1},\cdots,\lambda_{n})$ consisting of the eigenvalues. By a generalized Cauchy–Binet formula for permanents [MM65, Theorem 2.1], it holds that
@@
To relate the permanent view (29) with the doubly centered expansion in Section˜4, we refer to Appendix˜A for some identities between $S_{\ell}$ and related quantities. In particular, we note that the permanent $\mathrm{Perm}(U_{n-\ell,\ell,0,\ldots,0})$ is exactly a multiple of $e_{\ell}(u_{2})$, the elementary symmetric polynomial of second column vector $u_{2}$ of $U$. Therefore, Theorem˜4.3 could be used to upper bound this quantity. In addition, Lemma˜4.2 reduces the general quantity $\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})$ to the case $\mathrm{Perm}(U_{n-\ell,\ell,0,\ldots,0})$ with identical columns, so the arguments in Section˜4 leads to an upper bound of (29) in terms of $\sum_{i=2}^{n}\lambda_{i}\leq{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ by Lemma˜5.2. However, this approach does not capture additional structures of $(\lambda_{2},\dots,\lambda_{n})$ such as the spectral gap. To this end, it might be a natural idea to establish pointwise upper bounds on $\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})$. We have not succeeded in this approach and leave it as an open direction; instead, we choose to upper bound the individual sum $S_{\ell}$ or the entire sum $\sum_{\ell=0}^{n}S_{\ell}$. These form the topics of the subsequent sections.
5.2 Proof of Theorem˜1.2 via bounding the entire sum
In this section, we upper bound the entire sum $\sum_{\ell=0}^{n}S_{\ell}$ in (29) in order to prove Theorem˜1.2. The central result of this section is the following upper bound:
@@
Comparing this expression with (29), Lemma˜5.3 shows that replacing all coefficients in (29) by $1$ gives an upper bound. Using the eigenvalue properties in Lemma˜5.2, this upper bound implies the upper bounds (5) and (6) of Theorem˜1.2; see Section˜C.1. The proof of Lemma˜5.3 relies on the use of complex normal random variables. Recall that $z\sim{\mathcal{C}}{\mathcal{N}}(0,1)$ denotes $z=x+\mathrm{i}y$ with independent $x,y\sim{\mathcal{N}}(0,\frac{1}{2})$, and $z\sim{\mathcal{C}}{\mathcal{N}}(0,\Sigma)$ denotes a vector $z=x+\mathrm{i}y$ with independent $x,y\sim{\mathcal{N}}(0,\frac{\Sigma}{2})$ for real PSD matrices $\Sigma$. It is known that the moments of $z\sim{\mathcal{C}}{\mathcal{N}}(0,1)$ are
$\mathbb{E}[z^{m}\bar{z}^{n}]=n!\mathbbm{1}_{m=n}.$ (30)
@@
Lemma˜5.4 follows from the classical Isserlis’ theorem [Iss18] (or Wick’s formula) for products of joint normal random variables, and we include a proof in Section˜C.12 for completeness. Now we present the proof of Lemma˜5.3.
Proof of Lemma˜5.3.
Let $A=UDU^{\top}$ be the matrix in Lemma˜5.1, and $P=UD^{1/2}$. Then
@@
$\overset{\rm(a)}{=}\frac{1}{n^{n}}\sum_{\begin{subarray}{c}(\ell_{1},\ldots,\ell_{n})\in\mathbb{N}^{n}\\ \sum_{i=1}^{n}\ell_{i}=n\end{subarray}}\binom{n}{\ell_{1},\ldots,\ell_{n}}\lambda_{1}^{\ell_{1}}\cdots\lambda_{n}^{\ell_{n}}\mathbb{E}\left[|z_{1}|^{2\ell_{1}}\cdots|z_{n}|^{2\ell_{n}}\right]$
$\overset{\rm(b)}{=}\frac{n!}{n^{n}}\sum_{\begin{subarray}{c}(\ell_{1},\ldots,\ell_{n})\in\mathbb{N}^{n}\\ \sum_{i=1}^{n}\ell_{i}=n\end{subarray}}\lambda_{1}^{\ell_{1}}\cdots\lambda_{n}^{\ell_{n}},$
@@
5.3 Proof of Theorem˜1.5 via bounding the individual sum
In this section, we upper bound the individual sum $S_{\ell}$ in (29) in order to prove Theorem˜1.5. The central result of this section is the following upper bound: